OpenStudy (anonymous):

Help please! Write an equation in slope-intercept, point-slope or standard for the line with the given information. Explainh why you chose the for you used. a. passes thorugh (-1,4) and (-5,2)

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

do you know the forms for the equation of a line?

OpenStudy (anonymous):

nope thats all it gives me

OpenStudy (anonymous):

NOPE me & @batmanismyhero12 dont know the answers we both go to the same school

OpenStudy (anonymous):

ok so i will help you guys. from the question. you have 2 points. if you have 2 points. it means we can find the slope. right?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

ok. the slope of a line is given as say m. where m = (change in y) / (change in x). so for the given points. what will be our value for m?

OpenStudy (anonymous):

Given (x1,y1)(x2,y2) Find: \[m=\frac{ y2 - y1 }{ x2 - x1 }\] \[y - y1 = m(x - x1)\]

OpenStudy (anonymous):

@batmanismyhero12 try to solve for m, i will help u if u get stuck

OpenStudy (anonymous):

umm o god my brain is a a complete blank

OpenStudy (anonymous):

ok what are the values for x1 and x2 ?

OpenStudy (anonymous):

-1and -5

OpenStudy (anonymous):

y1 and y2?

OpenStudy (anonymous):

4 and 2

OpenStudy (anonymous):

ok what is y2-y1=?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2-4=?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

what is x2-x1?

OpenStudy (anonymous):

cool now plug into \[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

@Tiffany2012 wanna try it?

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

u got the numbers for y2-y1=-2 x2-x1=-4

OpenStudy (anonymous):

idk